# ChunkKV: Semantic-Preserving KV Cache Compression for Efficient Long-Context LLM Inference

### v1

@@ 4.1 In-Context Learning
Table 3 presents the performance comparison. The results demonstrate that ChunkKV outperforms other KV cache compression methods on different models and compression ratios. Table 4 presents the performance comparison of many-shot GSM8K, also ChunkKV outperforms other KV cache compression methods. The consistent superior performance of ChunkKV in both models underscores its effectiveness in maintaining crucial contextual information for complex arithmetic reasoning tasks.
Jailbreak In this section, we evaluate the performance of ChunkKV on the JailbreakV benchmark (Luo et al., 2024). The prompt settings are the same as those used by Luo et al. (2024).
Table 5 presents the performance comparison. The results demonstrate that ChunkKV outperforms other KV cache compression methods on different models and compression ratios. Which shows the effectiveness of ChunkKV in maintaining crucial contextual information for safety benchmark.
@@
Ratio StreamingLLM H2O SnapKV PyramidKV ChunkKV (Ours)
@@ 4.2 Long-Context Benchmark
LongBench and NIAH are two widely used benchmarks for KV cache compression methods. Both benchmarks have a context length that exceeds $10K$. NIAH requires retrieval capability, while LongBench is a meticulously designed benchmark suite that tests the capabilities of language models in handling extended documents and complex information sequences.
LongBench We use LongBench (Bai et al., 2024) to assess the performance of ChunkKV on tasks involving long-context inputs. For more details on LongBench, please refer to the APPENDIX F. We evaluated multiple KV cache eviction methods using the LongBench benchmark with LLaMA-3-8B-Instruct (Meta, 2024), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), and Qwen2-7B-Instruct (Yang et al., 2024a), with a KV cache compression ratio of $10\%$. The LongBench provides the Chinese subtask, and Qwen2-7B-Instruct also supports Chinese, so we tested Qwen2-7B-Instruct with different KV cache compression methods on the Chinese subtasks.
Tables 6 show that ChunkKV is capable of achieving on-par performance or even better than the full KV cache with less GPU memory consumption. This table presents the performance gap (in percentage) between each method and the FullKV baseline, where negative values indicate performance degradation compared to FullKV. The table is evaluated in the LongBench English subtask, where ChunkKV outperforms other compression methods overall. This suggests that ChunkKV’s approach of retaining semantic chunks is more effective in preserving important information compared to other discrete token-based compression methods. For detailed results and Chinese subtask results, please refer to Appendix B.2 and B.5.
@@
Ratio StreamingLLM H2O SnapKV PyramidKV ChunkKV (Ours)
@@
Needle-In-A-HayStack We use Needle-In-A-HayStack (NIAH) (Kamradt, 2023) to evaluate LLMs’ long-context retrieval capability. NIAH assesses how well LLM extract hidden tricked information from extensive documents, and follow LLM-as-a-Judge (Zheng et al., 2023) we apply GPT-4o-mini (OpenAI, 2023) to assess the accuracy of the retrieved information. We evaluated multiple KV cache eviction methods using NIAH with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2, setting benchmark context lengths to 8k and 32k tokens.
Table 7 provides statistical results for different compression methods. These findings clearly indicate the effectiveness of ChunkKV in managing varying token lengths and depth percentages, making it a robust choice for KV cache management in LLMs. Figure 3 presents the NIAH benchmark results for LLaMA-3-8B-Instruct. The vertical axis represents the depth percentage, while the horizontal axis represents the token length, with shorter lengths on the left and longer lengths on the right. A cell highlighted in green indicates that the method can retrieve the needle at that length and depth percentage. The detail visualization of the NIAH benchmark can be found in Appendix B.3. The visualization results demonstrate that ChunkKV outperforms other KV cache compression methods.
@@
KV cache Size StreamingLLM H2O SnapKV PyramidKV ChunkKV (Ours)
@@ 4.3 Index Reuse
The results in Table 8 shows that the layer-wise index reuse strategy (ChunkKV_reuse) further boosts performance, achieving up to a 20.7% reduction in latency, and throughput improvements are particularly notable for longer input sequences, with ChunkKV_reuse delivering up to a 26.5% improvement over FullKV.
@@
Figure 4: Comparison with different index reuse layers on LongBench.
Measuring Task Performance. This experiment evaluates the performance of the layer-wise index reuse approach by measuring the performance of the LongBench (Bai et al., 2024), the experiment settings are the same as LongBench in 4.2. And the number of index reuse layers is set from 1 to the number of layers in the model, where an index reuse layer of 1 corresponds to the normal ChunkKV without index reuse, and our method set reuse layer to 2.
Figure 4 illustrates the performance of ChunkKV with varying index reuse layers on the LongBench benchmark. Generally, reuse layer set to 2 can achieve the minimal performance degradation across all models. For more experiments on index reuse, please refer to the APPENDIX B.1.3.
@@ 5.1 Chunk Size
This section aims to investigate the impact of chunk size on the performance of ChunkKV. Different chunk sizes will lead to varying degrees of compression on the semantic information of the data. We set the experiemnt setting the same as in LongBench in Section 4.2. The chunk size is set from the range $\{1,3,5,10,20,30\}$. Figure 5 shows the performance of the ChunkKV with different chunk size on the LongBench and NIAH benchmarks. The three colorful curves represent three LLMs with different chunk sizes, and the colorful dashed line is the corresponding FullKV performance. For more experiments on the size of the chunks with different compression ratios, refer to the Appendix B.4.
Compression Chunk Size
Rate 1 3 5 10 15 20 30
10% 37.32 40.49 40.47 40.51 40.21 40.05 39.57
20% 38.80 40.66 40.57 40.74 40.53 40.46 40.04
30% 39.23 41.02 41.29 41.59 41.38 41.33 41.02
From Figure 5, we can observe that the LongBench performance of ChunkKV is not significantly affected by the chunk size, with performance variations less than 1%. The three curves are closely aligned, indicating that chunk sizes in the range of $\{10,20\}$ exhibit better performance.
Table 9 and 10 show the performance of ChunkKV with different comperession ratios and different chunk sizes on the LongBench and NIAH. We conducted extensive experiments across different compression ratios and KV cache sizes to shows the effectiveness of ChunkKV and the chunk size is robust.
KV Cache Chunk Size
Size 1 3 5 10 15 20 30
96 41.0 63.2 65.2 70.3 67.2 65.3 53.1
128 47.9 65.6 69.1 73.8 72.3 72.0 71.2
256 61.7 70.3 71.2 74.1 73.2 72.3 71.1
512 68.6 72.6 72.5 74.5 74.3 74.0 72.6
From the chunk size ablation study, we can observe that across different tasks (LongBench and NIAH) and various compression settings, a chunk size of 10 consistently delivers optimal or near-optimal performance. This empirical finding suggests that a chunk size of 10 strikes a good balance between preserving semantic information and compression efficiency, making it a robust default choice for ChunkKV. Therefore, we adopt this chunk size setting throughout our experiments.
@@ 6 Conclusion
We introduced ChunkKV, a novel KV cache compression method that preserves semantic information by retaining more informative chunks. Through extensive experiments across multiple state-of-the-art LLMs (including DeepSeek-R1, LLaMA-3, Qwen2, and Mistral) and diverse benchmarks (GSM8K, LongBench, NIAH, and JailbreakV), we demonstrate that ChunkKV consistently outperforms existing methods while using only a fraction of the memory. Our comprehensive analysis shows that ChunkKV’s chunk-based approach maintains crucial contextual information, leading to superior performance in complex reasoning tasks, long-context understanding, and safety evaluations. The method’s effectiveness is particularly evident in challenging scenarios like many-shot GSM8K and multi-document QA tasks, where semantic coherence is crucial. Furthermore, our proposed layer-wise index reuse technique provides significant computational efficiency gains with minimal performance impact, achieving up to 20.7% latency reduction and 26.5% throughput improvement. These findings, supported by detailed quantitative analysis and ablation studies, establish ChunkKV as a significant advancement in KV cache compression technology, offering an effective solution for deploying LLMs in resource-constrained environments while maintaining high-quality outputs.
Impact Statement
@@ A.2 Hypothetical Scenario
Both ChunkKV and discrete token-based methods would use this question to calculate observation scores for the document. However, their approaches to selecting and retaining information differ significantly.
@@ B.1.1 Efficiency Analysis
B.1.2 Layer-Wise Index Similarity
@@
B.1.3 Index Reuse Performance
Figure 16 illustrates the performance of ChunkKV with varying index reuse layers on the GSM8K benchmark. The experiment reveals that math problems are more sensitive to index reuse layers compared to LongBench. Both LLaMA3-8B-Instruct and Qwen2-7B-Instruct exhibit significant performance degradation, with LLaMA3-8B-Instruct experiencing a steeper decline after two layers of index reuse than Qwen2-7B-Instruct. This suggests that the Qwen2-7B-Instruct model may be more robust to index reuse.
Table 13 shows the performance of ChunkKV with different numbers of index reuse layers in GSM8K. The number of index reuse layers is set from 1 to the number of layers in the model, where a index reuse layer of 1 corresponds to the normal ChunkKV without index reuse, and 28/32 is the maximum number of layers for LLaMA-3-8B-Instruct and Qwen2-7B-Instruct. The significant performance drop of LLaMA-3-8B-Instruct raises another question: whether the KV cache compression method is more sensitive to the model’s mathematical reasoning ability.
Table 13: Reusing Indexing Performance Comparison on GSM8K
Model Number of Index Reuse Layers
1 2 3 5 8 10 20 28/32
LLaMA-3-8B-Instruct 74.5 74.6 65.9 44.1 15.3 2.20 1.60 1.80
Qwen2-7B-Instruct 71.2 71.2 73.0 69.4 67.4 71.1 54.0 49.4

### v3

@@ 4.1 In-Context Learning
GSM8K. In the in-context learning scenario, we evaluated multiple KV cache compression methods for GSM8K [23], which contains more than 1,000 arithmetic questions on LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct [26], Qwen2-7B-Instruct [28] and DeepSeek-R1-Distill-Llama-8B [25]. Following Agarwal et al. [34], we consider many-shot GSM8K as a long-context reasoning scenario, which is a more challenging task than long-context retrieval benchmark LongBench [21]. The CoT prompt settings for this experiment are the same as those used by Wei et al. [30], for many-shot GSM8K we set the number of shots to 50, where the prompt length is more than 4k tokens. For more details on the prompt settings, please refer to the APPENDIX G.
Table 3 presents the performance comparison. The results show that ChunkKV outperforms other KV cache compression methods on different models and compression ratios. Table 4 presents the performance comparison of many-shot GSM8K, and also ChunkKV outperforms other KV cache compression methods. The consistent superior performance of ChunkKV in both models underscores its effectiveness in maintaining crucial contextual information for complex arithmetic reasoning tasks by chunk level KV cache rather than the discrete token level.
@@
Ratio SLM H2O SKV PKV ChunkKV (Ours)
@@
Jailbreak. In this section, we evaluate the performance of ChunkKV on the JailbreakV benchmark [24], which is a safety Jailbreak benchmark for language models. The prompt settings are the same as those used by Luo et al. [24].
Table 5 presents the performance comparison. The results demonstrate that ChunkKV outperforms other KV cache compression methods on different compression ratios. This shows that for safety benchmark, the chunk level KV cache is more effective than other discrete token level compression methods.
@@ 4.2 Long-Context Benchmark
LongBench and NIAH are two widely used benchmarks for KV cache compression methods. Both benchmarks have a context length that exceeds $10K$. NIAH requires retrieval capability, while LongBench is a meticulously designed benchmark suite that tests the capabilities of language models in handling extended documents and complex information sequences. For more details on LongBench, please refer to the APPENDIX F.
LongBench. We use LongBench [21] to assess the performance of ChunkKV on tasks involving long-context inputs. We evaluated multiple KV cache eviction methods using the LongBench benchmark with LLaMA-3-8B-Instruct [26], Mistral-7B-Instruct-v0.3 [27], and Qwen2-7B-Instruct [28], with a KV cache compression ratio of $10\%$. LongBench-ZH provides the Chinese subtask, and Qwen2-7B-Instruct also supports Chinese, so we tested Qwen2-7B-Instruct with different KV cache compression methods on the Chinese subtasks.
@@
Ratio SLM H2O SKV PKV ChunkKV (Ours)
@@
Qwen2-7B-Instruct on LongBench-ZH FullKV: 38.60 $\uparrow$
10% -15.95% -5.31% +0.18% -5.31% +2.20%
Tables 6 present the performance gap (in percentage) between each method and the FullKV baseline, where negative values indicate performance degradation compared to FullKV. The table is evaluated in both the LongBench English and Chinese subtasks, where ChunkKV outperforms other compression methods overall. This suggests that ChunkKV’s approach of retaining semantic chunks is more effective in preserving important information compared to other discrete token-based compression methods. For the 70B model and Chinese subtask results, please refer to Appendices B.2 and B.5.
@@
KV cache Size SLM H2O SKV PKV ChunkKV (Ours)
@@
Needle-In-A-HayStack. We use NIAH [22] to evaluate the long-context retrieval capability of LLMs. NIAH assesses how well LLM extracts hidden tricked information from extensive documents, and following LLM-as-a-Judge [35] we apply GPT-4o-mini [36] to assess the accuracy of the retrieved information. We evaluated multiple KV cache eviction methods using NIAH with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2, setting benchmark context lengths to 8k and 32k tokens.
@@
Table 7 provides statistical results for different compression methods. These findings clearly indicate the effectiveness of ChunkKV in managing varying token lengths and depth percentages, making it a robust choice for KV cache management in LLMs. Figure 3 presents the NIAH benchmark results for LLaMA-3-8B-Instruct. The vertical axis represents the depth percentage, while the horizontal axis represents the token length, with shorter lengths on the left and longer lengths on the right. A cell highlighted in green indicates that the method can retrieve the needle at that length and depth percentage. The detailed visualization of the NIAH benchmark can be found in the Appendix B.3. The visualization results demonstrate that ChunkKV outperforms other KV cache compression methods.
@@ 4.3 Index Reuse
The results in Table 8 show that the layer-wise index reuse strategy (ChunkKV_reuse) further boosts performance, achieving up to a 20.7% reduction in latency, and throughput improvements are particularly notable for longer input sequences, with ChunkKV_reuse delivering up to a 26.5% improvement over FullKV. For more detailed results, please refer to Appendix B.7.
Table 9: Reusing Indexing Performance Comparison on LongBench and GSM8K. $\triangle$ indicates the performance degradation of index reuse compared to the baseline.
Model ChunkKV
Baseline Index Reuse
LongBench
LLaMA-3-8B-Inst 40.51 40.27-0.59%
Mistral-7B-Inst 46.71 46.43-0.59%
Qwen2-7B-Inst 40.88 40.76-0.29%
GSM8K
LLaMA-3-8B-Inst 74.5 74.6+0.13%
Qwen2-7B-Inst 71.2 71.2+0.00%
Measuring Task Performance. We evaluate the effectiveness of our proposed layer-wise index reuse technique on LongBench [21] and GSM8K [23] benchmarks. For these experiments, we use the same configuration as our main LongBench experiments in Section 4.2, with index reuse applied to consecutive layers (reuse layers = 2).
As shown in Table 9, layer-wise index reuse maintains the performance of the models while reducing computational requirements. In LongBench, performance degradation is minimal (less than 0.6%) for both models, while GSM8K shows neutral or slightly positive effects. This validates that semantic chunks selected by ChunkKV remain consistently important across adjacent transformer layers, enabling efficient computation without sacrificing accuracy. An additional analysis of different reuse depths and their impact on throughput is provided in Appendix B.1.2.
@@ 4.4 Chunk Size
This section aims to investigate the impact of chunk size on the performance of ChunkKV. Different chunk sizes will lead to varying degrees of compression on the semantic information of the data. We used the same experiment setting as in LongBench and NIAH (Section 4.2). The chunk size is set from the range $\{3,5,10,20,30\}$ under the compression rate 10% for LongBench and the 128 KV cache size for NIAH. For more experiments with different compression ratios, see Appendix B.4.
Table 10: LongBench and NIAH Results with Chunk Size Ablation
Model Full KV Chunk KV SnapKV H2O
size=3 size=5 size=10 size=20 size=30
LongBench $\uparrow$
LLaMA-3-8B-Instruct 41.46 40.49 40.47 40.51 40.05 39.57 40.15 37.06
Mistral-7B-Instruct 48.08 46.45 46.51 46.71 46.42 45.98 46.38 43.61
NIAH $\uparrow$
LLaMA-3-8B-Instruct 74.6 65.6 69.1 73.8 72.0 71.2 58.9 47.9
Mistral-7B-Instruct 99.8 98.1 99.2 99.8 99.8 99.1 91.6 88.2
As shown in Table 10, the performance remains relatively stable when the chunk size is between 5 and 20, with the best results consistently achieved at chunk size 10. When the chunk size is too small (e.g., 3), the context is fragmented, leading to a slight drop in performance. Conversely, when the chunk size is too large (for example, 30), the semantic granularity becomes too coarse, and important fine-grained information may be lost, also resulting in performance degradation. This trend is consistent across both LongBench and NIAH benchmarks, as well as across different model architectures, indicating that the optimal chunk size is not highly sensitive to the specific task or model. We also provide a line graph in Appendix B.4 to visually illustrate this trend. Based on these findings, we recommend using a chunk size of 10 as a robust default for most applications. For users with specific requirements, chunk size can be further tuned, but our results suggest that moderate values (5-20) generally offer a good trade-off between semantic preservation and compression efficiency.
4.5 Comparing with KV Quantization
For comprehensively evaluate the effectiveness of ChunkKV, we performed experiments comparing ChunkKV with the KIVI quantization methods [37]. Although both approaches aim to optimize LLM inference, they operate on fundamentally different principles: Quantization reduces KV matrix precision, whereas our eviction method reduces KV matrix size. For more detail, see Appendix B.6.
From an implementation perspective, quantization methods require the full KV cache during prefilling to produce quantized representations, which are then used during decoding. In contrast, ChunkKV employs token removal prior to prefilling, enabling operation with a compressed cache throughout the entire inference process. This distinction creates different efficiency profiles, and each method offers unique advantages.
Due to the fact that KIVI requires an old Python version, the ChunkKV results are not aligned with Table 8. The efficiency results in Table 11 reveal ChunkKV’s significant advantages in latency-critical metrics. Although both methods substantially reduce cache size (ChunkKV to 10% and KIVI-2bits to 15.63%), ChunkKV delivers superior metrics for Time to First Token (TTFT) and Token Processing Time (TPOT). Most remarkably, ChunkKV achieves a 164.66s total generation time compared to KIVI’s 226.52s at 2-bit quantization, representing a 27.3% improvement in overall inference speed.
Table 11: Efficiency Results for ChunkKV and KIVI on LlaMa-3-8B-Instruct
Configuration Prompt Output Compression Prefilling Cache TTFT TPOT Total Gen.
Length Length Ratio / nbits Time(s) $\downarrow$ Size(GB) $\downarrow$ (s) $\downarrow$ (ms) $\downarrow$ Time(s) $\downarrow$
FullKV 8192 4096 - 1.5621 1.0000 1.6013 45.9421 184.2934
KIVI 8192 4096 2bits 1.4024 0.1563 1.4325 54.9561 226.5234
KIVI 8192 4096 4bits 1.3916 0.2813 1.4146 52.0510 214.5634
ChunkKV 8192 4096 10% 1.3653 0.1000 1.3914 39.8702 164.6600
@@ 5 Conclusion
In this paper, we first indicate that current KV cache methods lack the semantic information of the data, and then we proposed a novel KV cache compression method that preserves semantic information by retaining more informative chunks. Through extensive experiments across multiple state-of-the-art LLMs (including DeepSeek-R1, LLaMA-3, Qwen2, and Mistral) and diverse benchmarks (GSM8K, LongBench, NIAH, and JailbreakV), we demonstrate that ChunkKV consistently outperforms existing methods while using only a fraction of the memory. Our comprehensive analysis shows that ChunkKV’s chunk-based approach maintains crucial contextual information, leading to superior performance in complex reasoning tasks, long-context understanding, and safety evaluations. The method’s effectiveness is particularly evident in challenging scenarios like many-shot GSM8K and multi-document QA tasks, where semantic coherence is crucial. Furthermore, our proposed layer-wise index reuse technique provides significant computational efficiency gains with minimal performance impact, achieving up to 20.7% latency reduction and 26.5% throughput improvement. These findings, supported by detailed quantitative analysis and ablation studies, establish ChunkKV as a significant advancement in KV cache compression technology, offering an effective solution for deploying LLMs in resource-constrained environments while maintaining high-quality outputs.
@@ A.2 Hypothetical Scenario
Both ChunkKV and discrete token-based methods would use this question to calculate attention scores for the document. However, their approaches to selecting and retaining information differ significantly.
@@ A.4 Implications for Model Performance
A.5 Implementation Details
In our implementation of ChunkKV, we focus on maximizing computational efficiency to ensure practical deployment in real-world applications. In algorithm 3, we adopt several key optimization strategies to reduce both computational overhead and memory footprint during inference.
Algorithm 3 ChunkKV Implementation
1:Input: $\mathbf{Q}\in\mathbb{R}^{T_{q}\times d}$, $\mathbf{K}\in\mathbb{R}^{T_{k}\times d}$, $\mathbf{V}\in\mathbb{R}^{T_{v}\times d}$, observe window size $w$, chunk size $c$, compressed KV cache max length $L_{\text{max}}$
2:Output: Compressed KV cache $\mathbf{K}^{\prime}$, $\mathbf{V}^{\prime}$
3:$q_{\text{observe}}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}$triangleright Extract observe window queries
4:$\mathbf{A}\leftarrow q_{\text{observe}}\mathbf{K}^{T}$triangleright $\mathbb{R}^{w\times T_{k}}$ attention scores
5:$C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$triangleright Number of chunks
[…]
7:Vectorized Chunk Score Calculation:
8:for $i=0$ to $C-1$ do
9:     $\text{start}\leftarrow i\cdot c$
10:     $\text{end}\leftarrow\min((i+1)\cdot c,T_{k})$
[…]
12:end for
13:$k\leftarrow\min(\left\lfloor\frac{L_{\text{max}}-w}{c}\right\rfloor,C)$triangleright Limit k to available chunks
[…]
15:Build selection mask:
16:$\text{mask}\leftarrow\text{zeros}(T_{k})$triangleright Binary mask for token selection
17:for each $i$ in kept_chunks do
18:     $\text{start}\leftarrow i\cdot c$
19:     $\text{end}\leftarrow\min((i+1)\cdot c,T_{k})$
[…]
21:end for
22:Ensure recent context is preserved:
[…]
24:Efficient compression:
25:$\text{indices}\leftarrow\text{nonzero}(\text{mask})$triangleright Get indices of tokens to keep
26:$\mathbf{K}^{\prime}\leftarrow\text{index\_select}(\mathbf{K},\text{indices})$
27:$\mathbf{V}^{\prime}\leftarrow\text{index\_select}(\mathbf{V},\text{indices})$
28:Return $\mathbf{K}^{\prime}$, $\mathbf{V}^{\prime}$
The optimized implementation features several key improvements:
Vectorized Operations. We leverage batch matrix operations for attention score calculation instead of explicit loops, significantly reducing computational overhead. By computing chunk scores through optimized tensor operations (lines 7-11), we achieve substantial acceleration compared to token-by-token processing.
Memory Optimization. Our implementation employs a binary mask approach (lines 15-20) that avoids redundant memory allocations and reduces fragmentation. This single-pass token selection strategy consolidates both chunk preservation and recent window retention operations, leading to more efficient memory usage during inference.
Boundary Handling. We incorporate robust boundary checks throughout the algorithm (lines 10, 12, 18) to handle edge cases such as incomplete chunks or limited compressed length. This ensures stable performance across variable input lengths without requiring special case handling.
Single-pass Compression. Rather than performing multiple concatenation operations, our implementation builds a comprehensive selection mask that includes both semantically important chunks and recent context tokens. This approach (lines 21-25) minimizes memory copying operations and reduces computational overhead.
Our PyTorch implementation further extends these optimizations by leveraging GPU acceleration for all vector and matrix operations. The integration of CUDA kernels for critical operations like attention score calculation and top-k selection enables real-time performance even with long context windows.
@@ B.1.1 Efficiency Analysis
B.1.2 Index Reuse Performance
Figure 5 and Table 13 illustrates the performance of ChunkKV with varying index reuse layers on the GSM8K benchmark. Figure 6 and Table 14 illustrates the performance of ChunkKV with varying index reuse layers on the LongBench benchmark. The experiment reveals that math problems are more sensitive to index reuse layers compared to LongBench. Both LLaMA3-8B-Instruct and Qwen2-7B-Instruct exhibit significant performance degradation, with LLaMA3-8B-Instruct experiencing a steeper decline after two layers of index reuse than Qwen2-7B-Instruct. This suggests that the Qwen2-7B-Instruct model may be more robust to index reuse.
Table 13: Reusing Indexing Performance Comparison on GSM8K
Model Number of Index Reuse Layers
1 2 3 5 8 10 20 28/32
LLaMA-3-8B-Instruct 74.5 74.6 65.9 44.1 15.3 2.20 1.60 1.80
Qwen2-7B-Instruct 71.2 71.2 73.0 69.4 67.4 71.1 54.0 49.4
Figure 6: Comparison with different index reuse layers on LongBench.
Table 14: Reusing Indexing Performance Comparison on LongBench
Model Number of Index Reuse Layers
1 2 3 5 10 20 28/32
LLaMA-3-8B-Instruct 40.51 40.27 39.45 40.05 39.59 37.42 35.67
Mistral-7B-Instruct 46.71 46.43 46.48 45.90 44.36 41.84 39.56
Qwen2-7B-Instruct 40.88 40.76 40.91 41.58 40.19 37.42 36.79
B.1.3 Layer-Wise Index Similarity